OpenStudy (summersnow8):

Calculate the energy change per mole for the following reaction? The mass of (2/1) H is 2.01355 amu and the mass of (4/2) He is 4.00150 amu. A) 2.30 x 109 kJ/mol B) 9.00 x 109 kJ/mol C) 1.79 x 1011 kJ/mol D) 3.52 x 1012 kJ/mol E) 2.84 x 1013 kJ/mol

OpenStudy (summersnow8):

\[\left(\begin{matrix}2 \\ 1\end{matrix}\right) H + \left(\begin{matrix}2 \\ 1\end{matrix}\right) H --> \left(\begin{matrix}4 \\ 2\end{matrix}\right)He \]

OpenStudy (aaronq):

so i think you're supposed to find the mass defect (the difference between them and use \(E=mc^2\)

OpenStudy (summersnow8):

@arronq I understand I use that equation, but how would i set it up